OpenStudy (anonymous):

The scores on an economics examination are normally distributed with a mean of 73 and a standard deviation of 16. If the instructor assigns a grade of A to 12% of the class, what is the lowest score a student may have and still obtain an A?

OpenStudy (anonymous):

So first we find the \(z\) score. \[ z=\frac{x-\mu}{\sigma} =\frac{x-73}{16} \]

OpenStudy (anonymous):

Now \(12\%\) gets and A, so we want to go from \(0\%\) to \(100\%-12\%=88\%\). So look up \(z=\Phi ^{-1}(0.880)\) to get the standard score. Then find the actual score by solving for \(x\)\[ \Phi ^{-1}(0.880)=\frac{x-73}{16} \]

OpenStudy (anonymous):

Now \(\Phi^{-1}(0.88)\approx 1.174987\) if you look it up. So\[ x \approx 16(1.174987)+73\approx 91.799792 \]So you would want to get \(92\) points to be safe.